The length of a rectangle is
1 ft
more than twice the width, and the area of the rectangle is
28 ft2
. Find the dimensions of the rectangle.

Answer by Boreal(15235)

W=x
L=2x+1
Area=(2x+1)x=28
2x^2+x=28
2x^2+x-28=0
(2x-7)(x+4)=0
x=3.5,-4. Only 3.5 makes sense.
2x+1=8
That area is 28 ft^2
width is 3.5 feet and length is 8 feet.